OpenStudy (anonymous):

Simplify the radical expression by rationalizing the denominator 9 square root 50 Divided square root of 50 can help please

OpenStudy (anonymous):

Is this your question? \[\frac{ 9\sqrt{50} }{ \sqrt{50} }\]